# Pre-trained Encoder Inference: Revealing Upstream Encoders In Downstream Machine Learning Services

### v1

@@ IV-A General Framework
IV-B Discussions
Hyperparameters selection. There are three hyperparameters in the PEI attack framework, which are: (1) $N$ the number of candidate encoders, (2) $M_{1}$ the number of objective samples, and (3) $M_{2}$ the number of PEI attack samples (per candidate encoder and objective sample). Increasing $N$ will make the $z$-test in the inference stage of the PEI attack more robust while increasing $M_{1}$ and $M_{2}$ will make the calculated PEI scores more reliable. In our experiments (Sections VI, VII, and VIII), $N$ is set as $6$ or $7$. We also empirically find that only setting $M_{1}$ to be $10$ (for image encoders) or $20$ (for text encoders) and $M_{2}$ to be $20$ is enough to synthesize effective PEI attack samples to launch the attack.
Query budget. We here focus on analyzing the black-box query budget of synthesizing PEI attack samples for single candidates. According to Algorithm 1, such a query budget for a single candidate encoder is $\mathcal{O}(M_{1}\cdot M_{2}\cdot C)$, where $\mathcal{O}(C)$ is the query budget for solving Eq. (1), which will be discussed in Section V. Combining with the aforementioned practical settings of $M_{1}$ and $M_{2}$ leads to a query budget of at most $\mathcal{O}(400\cdot C)$. In Sections VI-B and VII-B, we will further show that such a query budget is cheap in terms of price.
Comparison with existing works. We acknowledge that the idea of synthesizing data of similar embeddings has also been adopted by Lv et al. [22] to design encoder watermarks that can be verified in downstream services. However, their method is only a watermarking method, while the PEI is a type of privacy attack against pre-trained encoders. It is also worth noting that Lv et al. [22] is a white-box protection that requires accessing full parameters of pre-trained encoders to inject watermarks, while our PEI attack is a black-box attack that can work as long as black-box accesses to candidate encoders and targeted downstream services are provided.
V PEI Attacks on Different Data-modalities
This section presents detailed designs of the PEI attack against pre-trained encoders on different data modalities. Concretely, we will discuss how to solve Eq. (1) in Section IV-A via zero-order optimization to synthesize PEI attack samples for image encoders and text encoders.
Algorithm 2 Black-box PEI Attack Sample Synthesis (via Solving Eq. (1)) for Image Encoder
1:Candidate image encoder $f_{i}$, objective image sample $x^{(obj)}_{j}$, random sampling number $S$, perturbation radius $\epsilon>0$, training iteration $T$, learning rate $\eta$.
@@ V PEI Attacks on Different Data-modalities
8:     Update synthesized sample with gradient normalization:
@@
9:     $\quad x^{(atk)}_{i,j,k}\leftarrow x^{(atk)}_{i,j,k}-\frac{\eta\cdot\nabla_{x}\mathcal{L}_{i,j}(x^{(atk)}_{i,j,k})}{\|\nabla_{x}\mathcal{L}_{i,j}(x^{(atk)}_{i,j,k})\|_{\infty}}$.
10:     Clip $x^{(atk)}_{i,j,k}$ into the range $[0,1]^{\mathrm{dim}(\mathcal{X})}$.
11:end for
12:return $x^{(atk)}_{i,j,k}$
V-A Black-box Attack Samples Synthesis for Image Encoders
@@
When the candidate encoders $f_{1},\cdots,f_{N}$ are image encoders, the objective function $\mathcal{L}_{i,j}(x)$ in Eq. (1) can be seen as a continuous function concerning the input image $x$. Therefore, one can first leverage zeroth-order gradient estimation techniques [26, 27] to estimate the gradient of the objective function in a black-box manner and then use standard stochastic gradient optimization methods to solve Eq. (1). These techniques have also been widely adopted to attack black-box ML image models [50, 49].
@@
In this work, we use the two-point zeroth-order gradient estimation [27, 26] method to solve Eq. (1) for the image encoder $f_{i}$. Specifically, for the objective function $\mathcal{L}_{i,j}(x)$ in Eq. (1), its gradient $\nabla_{x}\mathcal{L}_{i,j}(x)$ can be estimated as
$\nabla_{x}\mathcal{L}_{i,j}(x)$
$\approx\frac{\dim(\mathcal{X})}{S}\sum_{s=1}^{S}\frac{\mathcal{L}_{i,j}(x+\epsilon\mu_{s})-\mathcal{L}_{i,j}(x-\epsilon\mu_{s})}{2\epsilon}\mu_{s},$ (6)
@@
where $S$ is the estimation random sampling number, $\epsilon>0$ is the estimation perturbation radius, and $\mu_{1},\cdots,\mu_{S}$ are i.i.d. random vectors in the same shape of $\mathcal{X}$ drawn from the unit sphere $\mathbb{S}^{\dim(\mathcal{X})-1}$. Here, a larger sampling number $S$ will result in a more accurate gradient estimation but a higher encoder querying budget, and vice versa. We further adopt $l_{\infty}$-norm gradient normalization during the optimization to relieve the difficulty of turning learning rate $\eta$ (since now we can explicitly specify the maximum pixel updating values). The overall procedures of solving Eq. (1) for image encoders are presented as Algorithm 2.
@@
Query budget. Combine Algorithm 2 with Algorithm 1, we have that the query budget of synthesizing PEI attack images is $\mathcal{O}(M_{1}\cdot M_{2}\cdot T\cdot 2S)$ per candidate encoder. Our experiments show that setting $M_{1},M_{2}$ to $10\sim 20$ and $T,S$ to $100\sim 200$ is enough to perform an effective PEI attack against image encoders. This results in an attack cost estimated based on real-world EaaS prices of only hundreds of dollars per candidate encoder (see Section VI-B).
V-B Black-box Attack Samples Synthesis for Text Encoders
When the candidate encoders $f_{1},\cdots,f_{N}$ are text encoders, minimizing the objective function $\mathcal{L}_{i,j}(x)$ in Eq. (1) becomes a discrete optimization problem since the input $x$ is discrete text data. This prevents one from applying the previous zeroth-order gradient estimation to text encoders.
To this end, we generalize the beam-search-based adversarial text synthesis method by Sadasivan et al. [25] from a white-box method to fit our black-box attack setting. Specifically, in our attack, the adversary will synthesize PEI attack texts of length $T>0$ character by character. During the whole synthesis process, at most $K_{1}>0$ candidate attack texts will be retained. In the $t$-th synthesis step, for each candidate text sample of length $(t-1)$, the adversary will concatenate it with $K_{2}>0$ randomly sampled characters to construct $K_{2}$ new samples of length $t$. This will result in $(K_{1}\times K_{2})$ new text samples of length $t$ at all. Then, the adversary will retain the $K_{1}$ samples from the $(K_{1}\times K_{2})$ ones with the smallest optimization losses (see Eq. (1)) to form the new candidate attack samples. The overall procedures of synthesizing PEI attack text samples via Eq. (1) are presented as Algorithm 3.
Comparison with Sadasivan et al. [25]. The original beam-search attack [25] is designed against white-box generative language models (GLMs) and thus could not be applied to our attacks against black-box text encoders. This is mainly due to the fact that the original method requires access to the tokenizer of the targeted models, which, however, is not known to the PEI adversary. We tackle this issue by introducing a new beam-search sampling strategy that performs uniform sampling on a pre-defined character set. Besides, it is also worth noting that currently, our new Algorithm 3 can only synthesize gibberish texts while Sadasivan et al. [25] can synthesize legible ones since it leverages the natural generative ability of the targeted GLMs. We believe by introducing additional GLMs, Algorithm 3 can be further improved to also synthesize legible PEI attack texts.
Query budget. Combine Algorithms 1 and 3, we know the query budget of synthesizing PEI attack texts is $\mathcal{O}(M_{1}\cdot M_{2}\cdot T\cdot K_{1}\cdot K_{2})$ per candidate. Our experiments show that setting $M_{1},M_{2}\leq 20$, $T\leq 32$, $K_{1}\leq 200$, and $K_{2}\leq 32$ is enough to perform an effective PEI attack against text encoders. This results in an estimated attack cost of hundreds of dollars per candidate encoder (see Sections VI-B and VII-B).
Algorithm 3 Black-box PEI Attack Sample Synthesis (via Solving Eq. (1)) for Text Encoder
1:Candidate text encoder $f_{i}$, objective text sample $x^{(obj)}_{j}$, a character set $C$, beam-search hyperparameters $K_{1}$ and $K_{2}$, PEI attack sample text length $T$.
2:synthesized PEI attack sample $x^{(atk)}_{i,j,k}$.
3:Initialize an empty array $arr$.
4:Insert an empty text sequence to $arr$.
5:Denote $\mathcal{L}_{i,j}(x):=\|f_{i}(x)-f_{i}(x^{(obj)}_{j})\|_{2}^{2}$.
6:for $t$ in $1,\cdots,T$ do
7:     Initialize two empty array $arr^{\prime}$ and $scores$.
8:     for $x$ in $arr$ do
9:         Uniformly randomly draw $K_{2}$ character $c_{1},\cdots,c_{K_{2}}$ from $C$ without replacement.
10:         for $s$ in $1,\cdots,K_{2}$ do
11:              $arr^{\prime}.\mathrm{append}(x\oplus c_{s})$triangleright $\oplus$ denotes concatenation.
12:              $scores.\mathrm{append}(\mathcal{L}_{i,j}(x\oplus c_{s}))$
13:         end for
14:         $arr\leftarrow\text{Top-}K_{1}(arr^{\prime},scores)$triangleright Select $K_{1}$ samples from $arr^{\prime}$ with lowest $scores$ to form a new $arr$
15:     end for
16:end for
17:Initialize an empty array $scores$.
18:for $x$ in $arr$ do
19:     $scores.\mathrm{append}(\mathcal{L}_{i,j}(x))$
20:end for
21:$x^{(atk)}_{i,j}\leftarrow\text{Top-}1(arr,scores)$triangleright Select the sample from $arr$ with the lowest $score$ as the eventual output.
22:return $x^{(atk)}_{i,j}$
VI Experiments on Image and Text Classifications
This section empirically analyzes PEI attack against image or text encoders hidden in image or text classification services.
@@ VI Experiments on Image and Text Classifications
TABLE I: PEI scores and PEI $z$-scores of candidate encoders on different downstream image classification services. If a candidate has $z$-score that is the only one above the threshold, then it is inferred as the encoder hidden in the downstream service.
Downstream Task Candidate Encoder PEI Score (%) / PEI $z$-Score (Threshold $=1.7$) Inferred Encoder Attack Success
@@
CIFAR-10 RN34 (HF) 74.50 / 2.04 3.00 / -0.33 0.00 / -0.43 0.00 / -0.43 1.00 / -0.40 0.00 / -0.43 RN34 (HF)
RN50 (HF) 11.50 / -0.46 57.50 / 1.98 10.50 / -0.52 22.00 / 0.09 10.50 / -0.52 9.50 / -0.57 RN50 (HF)
RN34 (MS) 39.50 / -0.38 40.00 / -0.34 78.00 / 2.03 36.50 / -0.56 39.50 / -0.38 39.50 / -0.38 RN34 (MS)
RN50 (MS) 2.00 / -0.56 15.50 / 0.03 0.50 / -0.62 60.50 / 1.99 3.50 / -0.49 7.00 / -0.34 RN50 (MS)
MobileNetV3 0.00 / -0.41 0.50 / -0.39 0.00 / -0.41 0.00 / -0.41 59.50 / 2.04 0.00 / -0.41 MobileNetV3
CLIP ViT-L/14 20.00 / -0.41 20.00 / -0.41 20.00 / -0.41 20.00 / -0.41 20.00 / -0.41 57.50 / 2.04 CLIP ViT-L/14
SVHN RN34 (HF) 75.00 / 2.03 15.50 / -0.41 16.00 / -0.39 15.00 / -0.43 20.50 / -0.20 11.00 / -0.59 RN34 (HF)
RN50 (HF) 28.50 / -0.52 64.00 / 2.03 30.50 / -0.38 32.00 / -0.27 30.50 / -0.38 29.00 / -0.49 RN50 (HF)
RN34 (MS) 23.00 / -0.07 15.00 / -0.58 55.00 / 1.95 15.00 / -0.58 24.50 / 0.02 12.50 / -0.74 RN34 (MS)
RN50 (MS) 20.50 / -0.21 19.00 / -0.40 21.50 / -0.08 35.50 / 1.74 24.00 / 0.25 12.00 / -1.31 RN50 (MS)
MobileNetV3 18.50 / -0.92 20.50 / -0.16 21.00 / 0.03 19.50 / -0.54 26.00 / 1.93 20.00 / -0.35 MobileNetV3
CLIP ViT-L/14 10.50 / -0.32 9.00 / -0.46 9.00 / -0.46 10.50 / -0.32 9.00 / -0.46 36.00 / 2.04 CLIP ViT-L/14
STL-10 RN34 (HF) 88.50 / 2.04 49.00 / -0.43 49.50 / -0.40 50.00 / -0.37 49.50 / -0.40 49.00 / -0.43 RN34 (HF)
RN50 (HF) 13.00 / -0.49 79.50 / 1.84 10.00 / -0.60 41.00 / 0.49 10.00 / -0.60 9.00 / -0.64 RN50 (HF)
RN34 (MS) 14.50 / -0.52 17.50 / -0.38 70.50 / 2.03 19.50 / -0.29 15.00 / -0.50 18.50 / -0.34 RN34 (MS)
RN50 (MS) 9.50 / -0.59 37.50 / 0.47 11.50 / -0.52 73.50 / 1.85 7.50 / -0.67 11.00 / -0.54 RN50 (MS)
MobileNetV3 2.00 / -0.43 3.00 / -0.40 3.00 / -0.40 5.50 / -0.31 72.00 / 2.04 0.00 / -0.50 MobileNetV3
CLIP ViT-L/14 6.00 / -0.50 8.50 / -0.39 6.00 / -0.50 9.50 / -0.35 11.00 / -0.29 65.00 / 2.03 CLIP ViT-L/14
Food-101 RN34 (HF) 43.50 / 2.04 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 RN34 (HF)
RN50 (HF) 1.00 / -0.38 27.50 / 2.04 0.50 / -0.43 1.50 / -0.33 0.50 / -0.43 0.00 / -0.47 RN50 (HF)
RN34 (MS) 0.00 / -0.41 0.00 / -0.41 25.00 / 2.04 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 RN34 (MS)
RN50 (MS) 0.50 / -0.36 0.00 / -0.44 0.00 / -0.44 15.00 / 2.04 0.50 / -0.36 0.00 / -0.44 RN50 (MS)
MobileNetV3 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 24.50 / 2.04 0.00 / -0.41 MobileNetV3
CLIP ViT-L/14 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 0.00 / -0.41 9.50 / 2.04 CLIP ViT-L/14
TABLE II: PEI scores and PEI $z$-scores of candidate encoders on different downstream text classification services. If a candidate has $z$-score that is the only one above the threshold, then it is inferred as the encoder hidden in the downstream service.
Downstream Task Candidate Encoder PEI Score (%) / PEI $z$-Score (Threshold = 1.7) Inferred Encoder Attack Success
Dataset Encoder BERT (base) BERT (large) T5 (small) T5 (base) RoBERTa (base) CLIP ViT-L/14
SST-5 BERT (base) 38.50 / 1.92 22.50 / -0.01 21.50 / -0.13 20.50 / -0.25 15.00 / -0.91 17.25 / -0.64 BERT (base)
BERT (large) 28.25 / -0.31 43.75 / 2.02 27.25 / -0.46 29.25 / -0.16 26.25 / -0.61 27.00 / -0.49 BERT (large)
T5 (small) 33.25 / 0.11 31.50 / -0.79 35.75 / 1.38 31.50 / -0.79 35.00 / 1.00 31.25 / -0.91 $\emptyset$
T5 (base) 38.50 / 0.14 38.75 / 0.21 35.25 / -0.77 38.50 / 0.14 33.25 / -1.33 43.75 / 1.60 $\emptyset$
RoBERTa (base) 24.50 / -0.63 19.75 / -1.53 28.00 / 0.04 29.00 / 0.23 35.25 / 1.42 30.25 / 0.47 $\emptyset$
CLIP ViT-L/14 27.75 / -0.57 25.25 / -0.88 33.50 / 0.14 32.00 / -0.05 28.00 / -0.54 47.75 / 1.90 CLIP ViT-L/14
TREC BERT (base) 63.75 / 2.00 30.00 / -0.39 27.00 / -0.60 30.25 / -0.37 35.00 / -0.04 27.00 / -0.60 BERT (base)
BERT (large) 35.25 / -0.13 50.25 / 1.35 43.75 / 0.71 33.25 / -0.32 36.50 / 0.00 20.25 / -1.60 $\emptyset$
T5 (small) 12.00 / -1.16 10.50 / -1.27 43.75 / 1.18 37.00 / 0.68 32.25 / 0.33 31.00 / 0.24 $\emptyset$
T5 (base) 22.75 / -1.33 29.00 / -0.47 34.00 / 0.21 45.00 / 1.71 32.75 / 0.04 31.25 / -0.17 T5 (base)
RoBERTa (base) 36.50 / -0.39 37.25 / -0.30 40.75 / 0.14 46.50 / 0.87 49.50 / 1.24 27.25 / -1.56 $\emptyset$
CLIP ViT-L/14 24.00 / -0.42 24.50 / -0.35 24.75 / -0.32 26.50 / -0.10 21.00 / -0.80 42.75 / 1.99 CLIP ViT-L/14
Yelp BERT (base) 46.50 / 1.80 31.25 / -0.73 30.00 / -0.94 32.75 / -0.48 36.00 / 0.06 37.50 / 0.30 BERT (base)
BERT (large) 38.50 / -1.63 55.50 / 1.20 52.75 / 0.74 45.00 / -0.55 48.25 / -0.01 49.75 / 0.24 $\emptyset$
T5 (small) 26.00 / -0.71 26.25 / -0.68 44.25 / 1.77 34.75 / 0.48 25.25 / -0.82 31.00 / -0.03 T5 (small)
T5 (base) 25.50 / 0.11 17.00 / -0.97 28.50 / 0.49 38.00 / 1.70 18.75 / -0.75 20.00 / -0.59 T5 (base)
RoBERTa (base) 26.00 / -0.88 26.25 / -0.81 29.75 / 0.16 27.75 / -0.39 35.75 / 1.84 29.50 / 0.09 RoBERTa (base)
CLIP ViT-L/14 13.50 / -0.94 20.00 / -0.18 19.75 / -0.21 21.25 / -0.03 16.75 / -0.56 38.00 / 1.93 CLIP ViT-L/14
AG-News BERT (base) 59.25 / 1.32 57.00 / 0.96 48.50 / -0.39 45.50 / -0.87 52.00 / 0.17 43.50 / -1.19 $\emptyset$
BERT (large) 27.75 / -0.37 49.25 / 1.98 31.75 / 0.06 27.00 / -0.46 26.00 / -0.57 25.25 / -0.65 BERT (large)
T5 (small) 54.75 / 0.97 54.00 / 0.73 54.75 / 0.97 47.75 / -1.26 49.25 / -0.78 49.75 / -0.62 $\emptyset$
T5 (base) 39.00 / -1.05 41.75 / -0.44 46.50 / 0.61 51.50 / 1.71 41.00 / -0.61 42.75 / -0.22 T5 (base)
RoBERTa (base) 84.75 / -0.77 86.25 / -0.13 89.25 / 1.17 85.50 / -0.45 89.50 / 1.27 84.00 / -1.10 $\emptyset$
CLIP ViT-L/14 40.50 / -0.77 45.00 / -0.13 47.00 / 0.15 39.50 / -0.92 44.50 / -0.20 59.00 / 1.87 CLIP ViT-L/14
VI-A Experimental Setup
Downstream datasets. For image classification, we adopt four image datasets as the downstream data, which are: CIFAR-10 [51], SVHN [52], STL-10 [53], and Food-101 [54]. For text classification, we adopt four text datasets as the downstream data, which are: SST-5 [55], Yelp [56], AG-News [56], and TREC [57]. Please refer to Appendix A-A for more details about these datasets.
Pre-trained encoders. For image classification, six image encoders are adopted as candidates in the PEI attack. They are: ResNet-34 (HF) [58], ResNet-50 (HF) [58], MobileNetV3 [59], ResNet-34 (MS) [58], ResNet-50 (MS) [58], and (the vision encoder of) CLIP ViT-L/14 [2]. The first three encoders are pre-trained by Hugging Face, the forth and fifth ones are pre-trained by Microsoft, and the last one is pre-trained by OpenAI. Besides, for text classification, we also leverage six text encoders as the candidates, which are: BERT (base) [1], BERT (large) [1], T5 (small) [5], T5 (base) [5], RoBERTa (base) [60], and (the language encoder of) CLIP ViT-L/14 [2]. The first four encoders are pre-trained by Google, the fifth one is pre-trained by Facebook, and the last one is pre-trained by OpenAI. These pre-trained image/text encoders can be downloaded from the Hugging Face Model Repository. See Table A-B2 in Appendix C for details.
Building downstream services. For each pair of encoder and downstream dataset, we fix the encoder and train a downstream classifier based on embeddings of downstream training data obtained from the encoder. For each service, the downstream classifier is an MLP consisting of three fully-connected layers, where the dimension of each hidden layer is $512$. We use Adam to train the classifier for $10,000$ iterations, where the batch size is set as $512$ and the learning rate is set as $0.001$ and decayed by a factor of $0.1$ every $4,000$ iterations.
PEI attack image synthesis. For image encoders, we randomly select $M_{1}=10$ images from the PASCAL VOC 2012 dataset [61] as the objective samples $x^{(obj)}_{1},\cdots,x^{(obj)}_{M_{1}}$. They are collected and presented as Figure VII in Appendix A-B. For each pair of candidate encoder $f_{i}$ and objective sample $x^{(obj)}_{j}$, we follow Algorithm 2 to synthesize $M_{2}=20$ PEI attack images $x^{(atk)}_{i,j,1},\cdots x^{(atk)}_{i,j,M_{2}}$. For the hyperparameters, without explicitly stating, the perturbation radius $\epsilon$ is set as $5.0$, the sampling number $S$ is set as $100$, the training iterations number $T$ is set as $200$, and the learning rate is fixed to $0.1$. The shape of each synthesized PEI attack image is $64\times 64$.
PEI attack text synthesis. For text encoders we manually construct $M_{1}=20$ sentences as the objective texts $x^{(obj)}_{1},\cdots,x^{(obj)}_{M_{1}}$, where each sentence follows the structure {Adjective} + {Noun}. They are collected in Table V in Appendix A-B. For each pair of candidate encoder $f_{i}$ and objective text $x^{(obj)}_{j}$, we follow Algoritm 3 to synthesize $M_{2}=20$ PEI attack texts $x^{(atk)}_{i,j,1},\cdots,x^{(atk)}_{i,j,M_{2}}$. For the hyperparameters, without explicitly stating, the character set is “abcdefghijklmnopqrstuvwxyz!?.,”, the text length $T$ is set as $16$, and the beam-search parameters $K_{1}$ and $K_{2}$ are set as $150$ and $16$ respectively.
Figure 3: PEI $z$-scores of candidate encoders on different CIFAR-10/SST-5 classification services where the correct hidden encoder is not in the PEI candidates set. $z$-scores that are above the preset threshold $1.7$ are highlighted. Ideally, none of the reported $z$-score should go beyond the preset threshold.
VI-B Results Analysis

### v2

@@ 3.2 General attack framework
Comparison with existing works. We acknowledge that the idea of synthesizing data of similar embeddings has also been adopted by lv2024ssl [46] to design encoder watermarks that can be verified in downstream services. However, their method is only a watermarking method, while the PEI is a type of privacy attack against pre-trained encoders. It is also worth noting that lv2024ssl [46] is a white-box protection that requires accessing full parameters of pre-trained encoders to inject watermarks, while our PEI attack is a black-box attack that can work as long as black-box accesses to candidate encoders and targeted downstream services are provided.
Algorithm 2 Black-box PEI Attack Sample Synthesis (via Solving Eq. (1)) for Vision Encoder
1:Candidate vision encoder $f_{i}$, objective image sample $x^{(obj)}_{j}$, random sampling number $S$, perturbation radius $\epsilon>0$, training iteration $T$, learning rate $\eta$.
@@
8:     Update $x^{(atk)}_{i,j,k}$ via gradient normalization: $\quad x^{(atk)}_{i,j,k}\leftarrow x^{(atk)}_{i,j,k}-\frac{\eta\cdot\nabla_{x}\mathcal{L}_{i,j}(x^{(atk)}_{i,j,k})}{\|\nabla_{x}\mathcal{L}_{i,j}(x^{(atk)}_{i,j,k})\|_{\infty}}$.
9:     Clip $x^{(atk)}_{i,j,k}$ into the range $[0,1]^{\mathrm{dim}(\mathcal{X})}$.
10:end for
11:return $x^{(atk)}_{i,j,k}$
3.3 PEI attack samples synthesis for vision encoders
@@
Finally, we present how to apply the proposed PEI attack (Algorithm 1) to threaten vision encoders deployed in downstream services. The main technical challenge lies in how to solve the PEI attack sample synthesis objective function $\mathcal{L}_{i,j}(x)$ defined in Eq. (1) in a black-box manner. Since when candidate encoders $f_{1},\cdots,f_{N}$ are vision encoders, the function $\mathcal{L}_{i,j}(x)$ can be seen as a continuous function for the input image $x$. Thus, one can leverage zeroth-order gradient estimation techniques to estimate the gradient of the function $\mathcal{L}_{i,j}(x)$ in a black-box manner and use gradient-based methods to solve Eq. (1). Such techniques are widely adopted to attack black-box vision models [36, 65].
@@
We use the two-point zeroth-order gradient estimation [44, 17] method to solve Eq. (1) for the vision encoder $f_{i}$. Specifically, for the objective function $\mathcal{L}_{i,j}(x)$ in Eq. (1), its gradient is estimated as
$\nabla_{x}\mathcal{L}_{i,j}(x)\approx\frac{\dim(\mathcal{X})}{S}\sum_{s=1}^{S}\frac{\mathcal{L}_{i,j}(x+\epsilon\mu_{s})-\mathcal{L}_{i,j}(x-\epsilon\mu_{s})}{2\epsilon}\mu_{s},$ (5)
@@
where $S$ is the estimation random sampling number, $\epsilon>0$ is the estimation perturbation radius, and $\{\mu_{s}\}_{s=1}^{S}$ are i.i.d. random vectors in the same shape of $\mathcal{X}$ drawn from the unit sphere $\mathbb{S}^{\dim(\mathcal{X})-1}$. A larger sampling number $S$ will result in a more accurate gradient estimation but a higher encoder querying budget. We further adopt $l_{\infty}$-norm gradient normalization during the optimization to relieve the difficulty of turning learning rate $\eta$ (as now we can explicitly control the maximum pixel updating values). The overall procedures of solving Eq. (1) for vision encoder are presented as Algorithm 2.
@@
Query budget. We focus on analyzing the black-box query budget of synthesizing PEI attack samples for single candidates. According to Algorithm 1, such a query budget for a single candidate encoder is $\mathcal{O}(M_{1}\cdot M_{2}\cdot C)$, where $\mathcal{O}(C)$ is the query budget for solving Eq. (1). Then, according to Algorithm 2, solving Eq. (1) for a single vision encoder requires a query budget of $\mathcal{O}(C)=\mathcal{O}(T\cdot 2S)$. All these lead to a query budget for synthesizing PEI attack images of $\mathcal{O}(M_{1}\cdot M_{2}\cdot T\cdot 2S)$ per candidate encoder. Our experiments in Section 4 and Section 5 show that setting $M_{1}=10$, $M_{2}=5$, $T=100$, and $S=100$ is enough to perform an effective PEI attack against vision encoders. This results in an attack cost estimated based on real-world EaaS prices of no more \$$\$100$ per candidate encoder.
4 Experiments on image classification services
In this section, we first analyze how the PEI attack can reveal vision encoders in image classification services and then show how revealed encoders can assist model stealing against targeted services.
4.1 Experimental setup
Building downstream services. We adopt three downstream image datasets, which are: CIFAR-10 [37], SVHN [47], and Food-101 [2]. Six vision encoders from the Hugging Face Model Repository are adopted as candidates in the PEI attack, which are: ResNet-34 (HF) [29], ResNet-50 (HF), MobileNetV3 [32], ResNet-34 (MS), ResNet-50 (MS), and CLIP ViT-L/14 [54]. The first three encoders were built by Hugging Face, the forth and fifth were built by Microsoft, and the last was built by OpenAI. For each pair of encoder and downstream dataset, we fix the encoder and train a downstream classifier, which is a MLP, based on embeddings of downstream training data obtained from the encoder. See Appendix B.1 for more details about the downstream services building.
PEI attack image synthesis. We randomly select $M_{1}=10$ images from the PASCAL VOC 2012 dataset [20] as the objective samples $\{x^{(obj)}_{i}\}_{i=1}^{M_{1}}$. They are collected and presented as Figure 4.3 in Appendix B.2. For each pair of candidate encoder $f_{i}$ and objective sample $x^{(obj)}_{j}$, we follow Algorithm 2 to synthesize $M_{2}=5$ PEI attack images $\{x^{(atk)}_{i,j,k}\}_{k=1}^{M_{2}}$. Meanwhile, the perturbation radius $\epsilon$ is set as $5.0$, the sampling number $S$ is set as $100$, the training iterations number $T$ is set as $100$, and the learning rate is fixed to $0.1$. The shape of each synthesized PEI attack image is $64\times 64$.
@@ 4.1 Experimental setup
Table 1: PEI scores and PEI $z$-scores of candidate encoders on different downstream image classification services. If a candidate has $z$-score that is the only one above the threshold, then it is inferred as the encoder hidden in the service. Blue color means that the hidden encoder is correctly revealed, while red color means that the PEI attack fails.
Downstream Task Candidate Encoder PEI Score (%) / PEI $z$-Score (Threshold $=1.7$) Inferred Encoder
@@
CIFAR-10 RN34 (HF) 52.0 / 2.04 2.0 / -0.33 0.0 / -0.43 0.0 / -0.43 0.0 / -0.43 0.0 / -0.43 RN34 (HF)
RN50 (HF) 14.0 / -0.28 50.0 / 2.01 10.0 / -0.53 16.0 / -0.15 10.0 / -0.53 10.0 / -0.53 RN50 (HF)
RN34 (MS) 36.0 / -0.82 40.0 / -0.20 54.0 / 1.94 38.0 / -0.51 42.0 / 0.10 38.0 / -0.51 RN34 (MS)
RN50 (MS) 0.0 / -0.77 22.0 / 0.47 2.0 / -0.65 46.0 / 1.81 2.0 / -0.65 10.0 / -0.21 RN50 (MS)
MobileNetV3 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 36.0 / 2.04 0.0 / -0.41 MobileNetV3
CLIP ViT-L/14 20.0 / -0.41 20.0 / -0.41 20.0 / -0.41 20.0 / -0.41 20.0 / -0.41 50.0 / 2.04 CLIP ViT-L/14
SVHN RN34 (HF) 58.0 / 1.98 4.0 / -0.86 14.0 / -0.33 16.0 / -0.23 14.0 / -0.33 16.0 / -0.23 RN34 (HF)
RN50 (HF) 30.0 / -0.28 46.0 / 1.99 30.0 / -0.28 30.0 / -0.28 30.0 / -0.28 26.0 / -0.85 RN50 (HF)
RN34 (MS) 16.0 / -0.40 16.0 / -0.40 36.0 / 1.78 16.0 / -0.40 24.0 / 0.47 10.0 / -1.06 RN34 (MS)
RN50 (MS) 24.0 / 0.70 12.0 / -0.98 16.0 / -0.42 30.0 / 1.54 20.0 / 0.14 12.0 / -0.98 $\emptyset$
MobileNetV3 18.0 / 0.00 20.0 / 0.32 24.0 / 0.97 20.0 / 0.32 6.0 / -1.94 20.0 / 0.32 $\emptyset$
CLIP ViT-L/14 10.0 / -0.24 10.0 / -0.24 10.0 / -0.24 10.0 / -0.24 8.0 / -0.98 16.0 / 1.95 CLIP ViT-L/14
Food-101 RN34 (HF) 32.0 / 2.04 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 RN34 (HF)
RN50 (HF) 0.0 / -0.67 16.0 / 1.79 0.0 / -0.67 8.0 / 0.56 0.0 / -0.67 2.0 / -0.36 RN50 (HF)
RN34 (MS) 0.0 / -0.41 0.0 / -0.41 6.0 / 2.04 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 RN34 (MS)
RN50 (MS) 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 4.0 / 2.04 0.0 / -0.41 0.0 / -0.41 RN50 (MS)
MobileNetV3 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 20.0 / 2.04 0.0 / -0.41 MobileNetV3
CLIP ViT-L/14 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 0.0 / -0.41 2.0 / 2.04 CLIP ViT-L/14
4.2 Results analysis